OpenStudy (anonymous):

find the of cot of 7pi/2 simple way

OpenStudy (dumbcow):

\[7\pi/2 = 2\pi + 3\pi/2\] \[\cot 3\pi/2 = undefined\]

OpenStudy (anonymous):

oh why is it 2pi + 3pi/2 shouldnt it be 3 and 1/2?

OpenStudy (dumbcow):

wait scratch that \[\tan 3\pi/2 = undefined\] so \[\cot 3\pi/2 = 0\]

OpenStudy (anonymous):

cot 7pi/2 =cos(630)/sin(630) =0

OpenStudy (anonymous):

im trying to use a method without calculator

OpenStudy (dumbcow):

use the unit circle \[\cot 3\pi/2 = \frac{\cos 3\pi/2}{\sin 3\pi/2} = 0\]

OpenStudy (anonymous):

pi=180 this means you multiply 7by 180 and divide the result by 2 cot=cos/sin meaning you 'ill have cos630/sin630 =0

OpenStudy (anonymous):

ok but can u explain how u got 7pi/2= 2pi + 3pi/2

OpenStudy (anonymous):

2pi + 3pi/2 = 4pi/2 + 3pi/2 = 7pi/2

OpenStudy (anonymous):

like i know 7/2= 3 1/2 but how does it add with 2pi to get 7/2

OpenStudy (anonymous):

I just showed u..

OpenStudy (anonymous):

oh ok looks bit confusing what is the purpose of those other fractions if u only need 3pi/2?

OpenStudy (anonymous):

I simply answered your question which was "how does it add with 2pi to get 7/2"

OpenStudy (anonymous):

better question is why do you have to show 2pi?

OpenStudy (anonymous):

tan and cot periodic (repeat values) in pi

OpenStudy (anonymous):

oh thanks everyone